The potential energy of a system of two atoms is given by the expression:

π‘ˆ = βˆ’ 𝐴 π‘Ÿ

2 ⁄ + 𝐡 π‘Ÿ

10 ⁄

A stable molecule is formed with release of 8.0 eV of energy when the interatomic distance isΒ 

2.8 Γ….Β 

i. Calculate A and B. [4 marks]

ii. Determine the force needed to dissociate this molecule into atoms and the interatomicΒ 

distance at which the dissociation occurs.Β 


1
Expert's answer
2021-01-19T12:22:57-0500

Answer

Energy is given

"E=8eV=12.8\times10^{-19}J"

Interatomic distance

"r=2.8A^°=2.8\times 10^{-10}m"

Potential energy is given

"U=-Ar^2+Br^{10}"

Now force ia given

"F=-\frac{dU}{dr}"

For equilibrium state

"F=0=-(-2Ar+10Br^9)"

By solving this

"r^8=\frac{A}{5B}"

"A=5Br^8"

"A=5\times (2.8\times10^{-10})^8B\\A=1.9\times10^{-76}B"

This relation shows value of A and B.


Dissociation energy is

"E=\frac{B^2}{4A}=\frac{B^2}{4\times1.9\times10^{-76}B}"

="8\times1.6\times10^{-19}" (Given)

After solving

B=51.2"\times10^{-95}" J/m^2

Now putting the value of B in above equation

"A=1.9\times10^{-76}\times 51.2\times10^{-95}"

"A=93.28\times10^{-171}J/m^{10}"